# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

### v1

@@ 4.4. Maximum Number of Vibration Units On Each Chain
Here we described how the maximum number of units on each chain was determined (DR1). Due to incremental wire resistance along the chains, voltage drops will become more significant when additional vibration units are connected or when more units are activated simultaneously. To maintain stable operation, the MCU power line must sustain a minimum of $2.3$ V (i.e., the MCU’s minimum operational voltage) and the actuator power line should not fall below $0.9$ V (i.e., the LRA-X actuator’s rated voltage).
Figure 11. Voltage drops on the MCU power line and the actuator power line with an increasing number of actuators being activated.
@@
We measured the voltage drop threshold by incrementally activating vibration units on the chain. An oscilloscope was used to measure the MCU power line and actuator power line at the last vibration unit of the chain. If the last unit maintained a stable voltage, then one could be certain that other units were stable as well. Figure 11 shows the relationship between voltage drop number of actuators activated. When the number of actuators surpassed 16, the MCU voltage dropped below the operating range. When it surpassed 17, the actuator voltage dropped below its rated voltage. Therefore, control units could support up to 16 actuators on each chain.
@@ 4.5. Vibration Characteristics
For intensity, we used an ADXL345 accelerometer for the measurements. Since most ADXL345 breakout boards are larger then the vibration unit itself, we designed a FlexPCB version and attached it between the skin and the vibration unit for precise measurements. A vibration unit was mounted on a sleeve and placed near the wrist of a human subject. During the testing, the vibration unit was activated for three seconds at each intensity level and the vibration frequency was kept constant at 170 Hz. The average acceleration was measured as the RMS value over the 1-second period of stable vibration. Acceleration was found to increase Monotonically as the intensity level increased (Figure 12 Left).
@@
Similarly, for frequency, the vibration unit was placed near the wrist and activated for three seconds at each frequency level with level-3 intensity (25% duty cycle). To obtain the main frequency response, the acceleration data was processed using a fast Fourier transform. We found that the main frequency response was well aligned with the designated frequency at each level, thereby demonstrating the precise PWM signal control (Figure 12 Right).
4.6. Waveform Approximation
Because vibration units are controlled via discrete UART commands rather than continuous audio signals, one might worry that this would reduce the expressivity of the vibration waveform delivered through the units. Here we show that by utilizing high-bandwidth communication (DR2), VibraForge was capable of approximating audio signals and creating similar feedback. We used standard vibration waveforms from VibViz (Seifi et al., 2015), converted them from WAV files to HAPTIC files through Meta Haptics Studio (Meta Platforms, 2023a), and sent them to the GUI Editor. The editor automatically converted the haptic files to a sequence of vibration commands that were used to drive a vibration unit. We then measured the acceleration of the vibration unit and compared it with the original audio envelope. The results showed that they were closely approximated to each other (Figure 13).
Figure 13. Comparison of actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library
@@ 4.6. Waveform Approximation
This figure compares actuator vibrations (green lines) and original audio waveforms (light green shaded areas) from the VibViz Library, showing four different waveforms.
@@ 5. Case Studies
We conducted three case studies to demonstrate the potential usage of VibraForge. The first case study showcased how VibraForge can be used to efficiently reproduce previous research prototypes. The second and third case studies showed how VibraForge can be used to build novel spatial vibrotactile systems that improve user performance in various applications.
@@ 5.1. Case Study 1: Replicating A Phonemic-Based Tactile Display
In Reed et al.’s work, a phonemic-based tactile display was designed to aid speech communication (Reed et al., 2018). This tactile display used 24 actuators on the forearm to deliver unique tactile codes of 39 English phonemes to deaf and hard-of-hearing users. Phonemes were mapped to vibration patterns using frequency, waveform, amplitude, spatial location, and movement. Although the display was shown to be useful, the process of creating the display was challenging. To drive the 24 actuators independently, the researchers had to use a desktop audio device (i.e., a MOTU 24Ao) with three custom amplifier boards to provide output signals, which made the whole system bulky and unable to be used in real-world situations. In this section, we demonstrated how the VibraForge toolkit could be utilized to simplify the development process and enhance the portability of the display.
5.1.1. Vibrotactile System Development
The first step in replicating the phonemic tactile display was to determine the layout of the actuators and the chain connections. Since the original display used a 4x6 layout on the forearm, we decided to use four chains with six vibration units on each chain and a small control unit.
@@ 5.1.1. Vibrotactile System Development
The top row shows all parts used for assembly, the assembly process itself, and the final prototype worn on an arm. The bottom row shows screenshots of waveform creation and layout design for the display, as well as an image of the original prototype used for comparison.
The second step was to assemble the display. The display required a sleeve garment and multiple modules from the toolkit (Figure 14). We first assembled the vibration units, then attached them to the garment and connected JST wires between them. The control unit was attached to the garment and four chains of vibration units were then connected to the control unit. The mapping between chains and connection ports followed the order of actuators described in the original paper. The whole assembly process took 16 minutes.
Once the hardware was built, the next step was to design the vibration patterns. Each phoneme needed to represented by a distinctive waveform that incorporated a high-frequency primary component and a low-frequency envelope, with a duration ranging from 100 to 480 ms. Consonants were designed as static patterns with multiple actuators being activated concurrently, whereas vowels were designed as dynamic patterns with actuators being activated in succession. The GUI Editor streamlined the process of creating the large variety of waveforms.
For example, to design the waveform for the consonant ”V” (IPA symbol /v/), we first dragged a sine wave into the Waveform Visualizer and set the frequency to be 300 Hz (primary component). Then we multiplied it with another sine wave at 8 Hz (envelope). The combined waveform was then saved in the Waveform Library as ”Consonant_B”. Alternatively, when the envelope was not available in the Waveform Library (e.g., $cos^{2}$ for consonant ”H”), we designed the waveform in Syntacts and imported the .CSV file into our GUI Editor. We then used the ”create new chain” function to create a virtual layout with four chains of vibration units on the canvas. We selected each of the designated vibration units (i.e., i6, ii6, iii6, iv6), assigned the custom waveform, and set the time range from 0 to 400 ms. Finally, we clicked the ”play” button to validate the design. The average time for an experienced haptic researcher on our research team to create one phoneme was one minute and five seconds. Therefore, the whole phoneme design process could be finished in 42 minutes. While the original paper did not mention the waveform design time, we believe the streamlined waveform creation process led to a significant reduction of design time.
@@ 5.1.2. Lessons Learned
The use of VibraForge to create the phonemic tactile display had several benefits. First, compared with the original device, our system was fully portable (DR5), with a small control unit and 24 vibration units directly attached to the sleeve. Second, VibraForge simplified the overall development process. The hardware system was reduced to module assembly without soldering, and the waveform design was supported by the GUI Editor (DR3). Third, the flexible design (DR4) enabled personalization opportunities. The original paper mentioned that the inter-motor spacing of the actuators needed to be adjusted for different participants and that the process of adjusting the layout was rather tedious. With our system, because the vibration units could be easily attached and detached, researchers could easily adjust the positions even if the garment was worn on the arm, thus giving them much more flexibility.
Meanwhile, some challenges were found when using the toolkit. While designers could easily author patterns using the GUI Editor, it mainly served as a testing site, because the current design of GUI Editor only allows users to manually click the ”play/stop” button to control the vibration sequences. This has constrained its use for real world interaction, where speech is segmented into phonemes and used to trigger corresponding patterns in real time. Right now the Python server does not support this kind of interaction so an additional API would be needed.
@@ 5.2. Case Study 2: Fitness Gaming in Virtual Reality
Virtual Reality (VR) fitness applications like OhShape (Lab, 2019) and Supernatural (Supernatural, 2024) have gained popularity by turning traditional fitness routines into interactive, gamified experiences. However, users often face challenges due to the absence of real-time, precise feedback about the physical positioning of their body (Wu et al., 2023). While prior research has demonstrated the utility of vibrotactile feedback for movement guidance in rehabilitation applications (Bark et al., 2014; Prabhu et al., 2020), systems used in these applications only deployed actuators on major joint locations such as the wrists or elbows, thus offering limited and coarse guidance. Inspired by OhShape (Lab, 2019) and the TV show ”Hole in the Wall” (Fox, 2008), we built a VR fitness application in which participants were asked to adjust their body poses to mimic cutouts in virtual walls when their digital avatar moved through the wall (Figure 15). We used the VibraForge toolkit to construct an upper-body suit that would provide spatial vibrotactile feedback to improve training performance.
Figure 15. An overview of the pipeline used in Case Study 2’s VR fitness game. Participant body movements were tracked by Vicon Motion Cameras and sent to a Unity scene for collision detection. Collision commands were then sent to a control unit on the participant’s back through Bluetooth and converted to UART commands for the vibration units.
@@
The Vicon motion capture cameras captured one’s motion and transmitted the body tracking data to a Unity scene, where it detected if a collision occurred between the participant and a virtual wall. If a collision occurred, then it activated the corresponding actuator to deliver tactile feedback.
@@ 5.2.1. Vibrotactile System Development
The first step of our development was to determine the layout of vibration units on the body. We referenced prior work that measured spatial acuity (Huang et al., 2024a) and determined that vibration units should be uniformly distributed on the arm with 4 cm of spacing between them and on the body with 8 cm of spacing between them. As a result, the suit had 36 vibration units on the body and 40 vibration units on each arm (Figure 15).
The second step was to mount the control units and vibration units on the garment, which was an exercise compression shirt. Since the total number of vibration units was within the upper limit of the large control unit, we only needed to use one control unit. It was worn on the upper back to help users maintain their balance during fitness training. Vibration units were mounted one-by-one and JST wires were used to connect them.
The third step was to integrate the hardware into the fitness game, which was built using the Unity Game Engine (Technologies, 2020). The physical layout of the vibration units on the suit was manually converted into a virtual layout and overlaid on the user’s avatar, so that the virtual vibration units would move following the human body movement. Whenever a collision was detected between a virtual vibration unit and an obstacle, the system would determine which vibration units should be activated and send the vibration commands to the control unit through a Python server (Figure 15). Since collisions were binary events (i.e., collided / not collided), the vibration intensity was set to constant level (i.e., 7), as well as frequency (i.e., 170 Hz).
@@ 5.2.2. User Study Design
A preliminary study was conducted to evaluate the effectiveness of the vibrotactile suit on fitness training performance. The VR application was run on an Oculus Quest 2 headset (Meta Platforms, 2023b) through a Quest Link connection to a PC (i.e., Alienware M15 Laptop with RTX 3060Ti GPU). Vicon Motion Capture cameras (Vicon, 2023) were used to track the body poses of each participant and the poses were then mapped to the virtual avatar in the Unity scene. Nine participants (i.e., 4 Males and 5 Females; mean age = 25 years, SD = 3 years) were recruited to participate in the study, which lasted approximately one and a half hours, and each received \$40 compensation. The study was approved by our institutional research ethics board and conducted in a university lab space.
During the study, each participant would experience two conditions, i.e., no feedback and vibrotactile feedback. In the no feedback condition, participants only saw the incoming virtual walls and their own avatar and had to rely on proprioception to adjust their body position. In the vibrotactile feedback condition, participants would feel vibrations at the body positions where collisions occurred between vibration units and walls. For each condition, 24 pairs of walls with different cutouts were presented in sequence. Two walls in each pair were identical. The first wall served as the ”preview wall” to help the participant adjust their body pose, and the second wall served as the ”test wall” to measure any body position changes. We measured the number of collisions occurring during each condition. Our hypotheses were that (1) vibrotactile feedback would cause fewer collisions than no feedback, and (2) test walls would have fewer collisions than preview walls.
@@ 5.2.4. Lessons Learned
The benefits of using VibraForge to build this high-density vibrotactile suit were multi-fold. First, scalability (DR1) enabled us to support a large number of actuators and provide highly localized feedback to guide body movement. Second, the modular and flexible design (DR4) made the system assembly process as simple as press-fitting the vibration units onto the garment and connecting wires between them. The entire process took less than 30 minutes. Compared with prior work that used sewing techniques (West et al., 2019; Lemmens et al., 2009), our solution was more flexible and efficient. Third, the low latency and high bandwidth communication (DR2) helped deliver almost real-time vibrotactile feedback to users when collisions occurred. Participants did not report any perceivable delays during the study.
Nonetheless, there were several challenges during the process. First, the integration of the vibrotactile system into the interactive game design was difficult. This was because we had to manually convert the physical layout of units on the body to a virtual layout in the Unity Engine and assign vibration units to their positions. This process would have been easier if we could detect the physical positions of the units and generate a virtual mapping automatically. Second, two participants reported that vibration units near the waist were not consistently perceived because body movement during fitness training would loosen the contact with garment. Our temporary solution was to tighten the suit so that movement was constrained. In the future, other methods would be needed to ensure that the vibration units maintained contact with the skin.
5.3. Case Study 3: Teleoperation Assistance for Unmanned Aerial Vehicle Operation
In this case study, we used VibraForge to design a wearable vibrotactile device to provide Unmanned Aerial Vehicle (UAV) operators with spatial information about obstacles. Current UAV teleoperation is dependent on visual feedback from onboard cameras, which presents significant challenges as a camera’s limited field of view can hinder situation awareness and lead to potential collisions with obstacles. While there have been prior attempts to use haptic feedback to notify operators of nearby obstacles (Brandt, 2009; Hou et al., 2017; Zhang et al., 2020; Xia et al., 2022), they required an operator to switch from using radio control (RC) controllers to hand-held haptic joysticks, only rendered force feedback in one direction at a time, or had fixed actuator layouts that could not represent lateral movements due to the absence of actuators on the side.
By creating a new vibrotactile feedback system using VibraForge, one could flexibly attach multiple vibration units to an operator’s body that would render simultaneous spatial cues about surrounding obstacles while the operator was using an RC controller. This would enhance operator’s collision avoidance ability and situational awareness in environments where visual feedback is limited.

### v2

@@ 4.4. Maximum Number of Vibration Units On Each Chain
Here we described how the maximum number of units on each chain was determined (DR1). Due to incremental wire resistance along the chains, voltage drops become more significant when additional vibration units are connected or activated simultaneously. To maintain stable operation, the MCU supply must maintain above $2.3$ V (i.e., the MCU’s minimum operational voltage) and the actuator supply should not fall below $0.9$ V (i.e., LRA-X actuator’s rated voltage).
Figure 11. Results of the voltage drops on the MCU (red) and the actuator (blue) power supply with an increasing number of activated actuators.
@@
We measured the voltage drop threshold by incrementally activating vibration units on a chain. An oscilloscope measured the MCU power line and actuator power line at the last vibration unit of the chain. If the last unit maintained a stable voltage, then one could assume the other units were stable as well. When the number of actuators surpassed 16, the MCU voltage dropped below the operating range (Figure 11). When it surpassed 17, the actuator voltage dropped below its rated voltage. Therefore, control units could support up to 16 actuators on each chain.
@@ 4.5. Vibration Characteristics
For intensity, we used designed a FlexPCB version of the ADXL345 accelerometer and attach it between the skin and the vibration unit. A vibration unit was mounted on a sleeve and placed near the wrist of a human subject. During testing, the vibration unit was activated for three seconds at each intensity level and the vibration frequency was kept constant at 170 Hz. The average acceleration was measured as the RMS value over the 1-second period of stable vibration. Acceleration was found to increase monotonically as intensity increased (Figure 12 Left).
Similarly, for frequency, the vibration unit was placed near the wrist and activated for three seconds at each frequency level with an intensity of 25% . The acceleration data was processed using a fast Fourier transform and the resulting main frequency responses were found well aligned with the designated frequency at each level, thereby demonstrating the precise PWM signal control (Figure 12 Right).
4.6. Waveform Approximation
Because vibration units are controlled via discrete UART commands rather than continuous audio signals, one might worry that this would reduce the expressivity of the vibration waveform delivered through the units. By utilizing high-bandwidth communication (DR2), VibraForge was capable of approximating audio signals and creating similar feedback. To demonstrate this, we used standard vibration waveforms from VibViz (Seifi et al., 2015), converted them from WAV files to HAPTIC files using Meta Haptics Studio (Meta Platforms, 2023a), and sent them to the GUI Editor for conversion into a sequence of vibration commands to drive a vibration unit. After measuring the acceleration of the vibration unit and comparing it to the original audio envelope, both measurements closely approximated each other (Figure 13).
@@ 4.6. Waveform Approximation
Figure 13. Comparison of actuator vibrations (blue lines) and original audio waveforms (light blue lines) from the VibViz Library.
@@
This figure compares actuator vibrations (blue lines) and original audio waveforms (light blue shaded areas) from the VibViz Library, showing four different waveforms.
4.7. Cost Estimation
VibraForge is a cost-effective toolkit. The small control unit costs \$27.23 USD (i.e., \$12.50 for the MCU, \$10.50 for the PCB, \$4.00 for the LiPo batteries, and \$0.23 for the 3D printed parts), which is similar to Arduino Uno but cheaper than Arduino Mega2560 and Raspberry Pi. The vibration unit costs \$3.46 USD (i.e., \$2.43 for the PCB, \$1.00 for the LRA, and \$0.03 for the 3D printed parts). For the wearable vibrotactile device in Case Study 1, for example, the total cost of a small control unit and 24 vibration units would be \$ 110.27 USD. This low manufacturing cost would enable designers to reproduce our toolkit with low financial burden. Note that these estimates are based on mass production (e.g., 10 control units and 100 vibration units), so prices may increase when manufacturing in smaller quantities.
@@ 5. Case Studies
We conducted three case studies to investigate the potential “ceiling” (Ledo et al., 2018; Myers et al., 2000) of the toolkit’s design space, measured by scalability and expressivity. The first case study demonstrated the toolkit’s capability for high expressivity by replicating a phonemic display that requires complex vibration patterns. The second case study showcased high scalability by utilizing a larger number of units for VR interaction. The third case study found a balance between the two metrics, where parameterized vibrations were used to convey collision risks during drone teleoperation.
@@ 5.1. Case Study 1: Replicating A Phonemic-Based Tactile Display
In Reed et al.’s work, a phonemic-based tactile display used 24 actuators on the forearm to deliver tactile codes of 39 English phonemes to deaf and hard-of-hearing users(Reed et al., 2018). Phonemes were mapped to vibration patterns using frequency, waveform, amplitude, spatial location, and movement. Although the display was found to be useful, the building process was challenging. A desktop audio device (MOTU 24Ao) with three custom amplifier boards was needed to drive the 24 actuators independently, which made the system bulky and unusable in real-world situations. In this section, we demonstrated how the VibraForge toolkit could simplify the development process, enhancing usability and portability.
@@
The top row shows the parts used for assembly, the assembly process itself, and the final prototype on an arm. The bottom row shows screenshots of waveform and layout creation, as well as an image of the original prototype for comparison.
5.1.1. Vibrotactile System Development
The first step was to determine the layout of the actuators and the chain connections. Since the original display used a 4x6 layout on the forearm, we decided to use four chains with six vibration units on each chain and a small control unit. Then we assembled the vibration units, attached them to the garment, and used JST wires to connect them (Figure 14). The control unit was then attached to the garment and four chains of vibration units were connected to the control unit. The mapping between chains and connection ports used the order of actuators described in the original paper. The assembly process took 16 minutes.
Once the hardware was built, we designed vibration patterns in the GUI Editor. Each phoneme was represented by a distinctive waveform that incorporated a high-frequency primary component and a low-frequency envelope, with a duration between 100 and 480 ms. Consonants were designed as static patterns with multiple actuators activated concurrently, whereas vowels were designed as dynamic patterns with actuators activated in succession.
To design the waveform for the consonant ”V” (IPA symbol /v/), for example, we began by dragging a 300 Hz sine wave into the Waveform Visualizer and multiplying it with an 8 Hz sine wave, saving the result as a custom waveform in the library. If an envelope waveform like $cos^{2}$ (for consonant ”H”) was unavailable, we created it in Syntacts (Pezent et al., 2020) and imported it into the Editor. We then used the ”create new chain” function to design a layout with four chains of vibration units on the canvas. We selected each of the designated units (i.e., i6, ii6, iii6, iv6), assigned the custom waveform, set the duration as 400 ms. Finally, we validated the design by clicking the ”Play” button. An experienced haptic researcher averaged one minute and five seconds to design a phoneme, and would thus complete the entire process in 42 minutes. This streamlined design process would foster rapid design iterations.
@@ 5.1.2. Lessons Learned
The use of VibraForge to create the phonemic tactile display offered several advantages. First, it was fully portable (DR5) with a compact control unit and 24 vibration units integrated into the sleeve. Second, VibraForge streamlined the development process with solder-free module assembly and rapid waveform design in the GUI Editor (DR3). Third, its flexible design (DR4) enable easy position adjustments, addressing the design challenge of personalization mentioned in the original paper.
Meanwhile, some challenges remained. While the GUI Editor allowed for easy pattern creation, it mainly served as a testing site and lacked integration into real applications. If the designer wanted to play vibration patterns when certain phonemes are detected in speeches, the designer needed to manually convert patterns into sequences of vibration commands to fit the APIs. New API functions are needed to support the integration of GUI-designed waveforms into external programming environment.
@@ 5.2. Case Study 2: Fitness Gaming in Virtual Reality
Virtual Reality (VR) fitness applications like OhShape (Lab, 2019) and Supernatural (Supernatural, 2024) have turned traditional fitness routines into interactive, gamified experiences. However, the absence of real-time and precise feedback about the physical positioning of one’s body makes them difficult to use (Wu et al., 2023). While prior work has demonstrated the utility of vibrotactile feedback for movement guidance (Bark et al., 2014; Prabhu et al., 2020; Jo et al., 2023), actuators were only deployed on major joint locations (e.g., wrists or elbows), thus offering limited and coarse guidance. Inspired by the TV show ”Hole in the Wall” (Fox, 2008), we built a VR fitness application that asked participants to adjust their body to mimic cutouts in virtual walls (Figure 15). VibraForge was used to construct an upper-body suit that provided localized vibrotactile feedback to improve user performance.
Figure 15. An overview of the pipeline used in Case Study 2, where participant body movements were tracked by Vicon Motion Cameras and sent to a Unity scene for collision detection. Collision commands were then sent to a control unit on the participant’s back through Bluetooth and converted to UART commands for the vibration units.
@@
The Vicon motion capture cameras captured one’s motion and transmitted the body tracking data to a Unity scene, where it detected if a collision occurred between the participant and a virtual wall and activated the corresponding actuator to deliver tactile feedback.
@@ 5.2.1. Vibrotactile System Development
The first step was to determine the suit layout and assemble the hardware components. We referenced prior work on spatial acuity (Elsayed et al., 2020; Huang et al., 2024a) and determined that vibration units should be uniformly distributed on the arm with a spacing of 4 cm and on the body with an 8 cm spacing. The suit thus had 36 vibration units on the body and 40 on each arm (Figure 15). The large control unit was used to drive the substantial number of vibration units, and a compression shirt was used as the garment. We affixed the control unit to the upper back to help maintain balance. Then we affixed the vibration units and connected them with JST wires.
Once the suit was assembled, we integrated the hardware into the fitness game, which was built using the Unity Engine (Technologies, 2020). The physical layout of the vibration units was manually converted into a virtual layout and overlaid on an avatar so that the virtual units would follow human body movement. Whenever a collision was detected between a virtual unit and a wall, the system sent vibration commands to the control unit via Unity plugin APIs and activated corresponding units (Figure 15). Since collisions were binary (i.e., collided / not collided), the vibration intensity and frequency were set constant (i.e., 50% and 170 Hz).
@@ 5.2.2. User Study Design
A preliminary study evaluated the effectiveness of the vibrotactile suit on fitness training performance. The VR application was run on an Oculus Quest 2 headset (Meta Platforms, 2023b) through a Quest Link connection to an Alienware M15 Laptop. Vicon Motion Capture cameras (Vicon, 2023) tracked the body pose of each participant and mapped it to a virtual avatar in the Unity scene. Nine participants (i.e., 4 Males and 5 Females; mean age = 25 years, SD = 3 years) were recruited for the 1.5-hour study, and each received \$40 compensation. The study was approved by our institutional research ethics board and conducted in a university lab space.
During the study, each participant would experience two conditions, i.e., no feedback and vibrotactile feedback. In the no feedback condition, participants only saw the incoming virtual walls and their own avatar and relied on proprioception to adjust their body position. In the vibrotactile feedback condition, participants would feel vibrations at the body positions where collisions occurred between vibration units and walls. For each condition, 24 pairs of walls with different cutouts were presented in sequence. Two walls in each pair were identical. The first wall served as the ”preview wall” to help the participant adjust their body pose, and the second wall served as the ”test wall” to measure any body position changes. We measured the number of collisions occurring during each condition. Our hypotheses were that (1) vibrotactile feedback would cause fewer collisions than no feedback, and (2) test walls would have fewer collisions than preview walls.
@@ 5.2.4. Lessons Learned
Using VibraFogre to build the high-density vibrotactile suit offered several advantages. First, scalability (DR1) supported numerous vibration units (116 in total), providing localized feedback for guiding body movements. Second, the modular and flexible design (DR4) reduced assembly to less than 30 minutes with simple press-fit attachment and rapid adjustments. Third, low latency and high bandwidth communication (DR2) ensured real-time feedback when collisions happened, with no perceivable delays reported.
Nonetheless, some challenges arose. The process of manually mapping the vibration units to the virtual layout in Unity was rather tedious. This could be solved by a localization pipeline that automatically detects vibration units’ positions and their relative orders. Additionally, two participants reported inconsistent vibrations near the waist due to loosened fabric contact and body movement. We addressed the problem by temporarily tightening the suit, but future solutions are needed to ensure consistent skin contact.
5.3. Case Study 3: Collision Avoidance for Unmanned Aerial Vehicle (UAV) Teleoperation
In UAV teleoperation tasks, operators heavily rely on visual feedback to detect obstacles and avoid collisions, but the narrow field of view from onboard cameras limits their situational awareness and increases collision risks. While prior work attempted to use haptic feedback for notifying operators of obstacles (Brandt and Colton, 2010; Hou et al., 2017; Zhang et al., 2020), they required operators to switch from radio control (RC) controllers to unfamiliar haptic joysticks, and provided only one directional force feedback. A spatialized vibrotactile feedback system using VibraForge could deliver simultaneous spatial cues about multiple obstacles, while operators continued using RC controllers (Figure 16). Additionally, vibration characteristics can be used to convey different levels of risks, helping operators prioritize high-risk obstacles.
Figure 16. The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point bodily vibrotactile feedback. Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange).
[…]
A human operator wearing the vibrotactile device is flying a simulated drone. He perceives incoming obstacles through vibrotactile feedback triggered on the body. He not only sees visible obstacles but also perceives obstacles outside the field of view.